Suppose that we have a system in which the energy levels are evenly spaced, and that when the system transitions from one state to the next lowest state a photon of wavelength 800 nm is emitted. What is the wavelength (in nm) if the system drops by two energy levels instead of one?
1
Expert's answer
2020-12-01T14:25:27-0500

Wavelength=800nm

Calculate the transaction energy of the 2levels.

Energy of the absorbed photons is calculated using the following formula;

"E=\dfrac{hc}{\lambda}"


"\dfrac{(6,63*10^{-34})*(3*10^8)}{800*10^{-9}}"


"=2,4*10^{-19}J"

If the system drops by 2 levels then

"=2,4*10^{-19}*2\newline =4.8*10^{-19}"

Wavelength ("\lambda") "=\dfrac{hc}{E}"


"=\dfrac{(6,63*10^{-34})*(3*10^8)}{4,8*10^{-19}J}"


"=414nm"